# Guided simulation of conditioned chemical reaction networks

### v1

@@ 4. Choices for $g$
What remains is a specification of the maps $(t,x)\mapsto g(t,x)$ satisfying the assumptions of 3.4. In this section, we demonstrate that the results in Fearnhead, (2008) and Golightly & Sherlock, (2019) can be considered as specific specific choices of $g$ satisfying these assumptions. We also propose a more easily computable choice of $g$ which also yields absolute continuity. Lastly we discuss a choice for $g$ that, under one additional assumption on the reaction network, yields equivalence of the measures ${\mathbb{P}}_{T}^{h}$ and ${\mathbb{P}}_{T}^{g}$ (rather than absolute continuity). In the following, $C$ denotes a covariance matrix.
@@ 4.1. Choices based on the chemical Langevin equation
$g_{\rm F}(t,x)={\mathcal{N}}\left(v;L(x+b_{\mathrm{CLE}}(t,x)(T-t)),La_{\mathrm{CLE}}(t,x)L^{\prime}(T-t)+C\right),$
Since $g$ is obtained as a composition of continuously differentiable maps, we infer absolute continuity $\mathbb{P}^{h}_{T}\ll\mathbb{P}^{g}_{T}$ from 3.6.
Golightly & Sherlock, (2019) proposed to infer $g$ by using the Linear Noise Approximation (LNA) to the CLE. Here, we discuss their specific choice called “LNA with restart”.
For notational convenience, write $b=b_{\mathrm{CLE}}$ and $\sigma=\sigma_{\mathrm{CLE}}$. Given $t<T$ and $x\in{\mathbb{S}}$, we denote by $z_{T\mid(t,x)}$ and $V_{T\mid(t,x)}$ the solutions at time $T$ of the system of ordinary differential equations
@@
Absolute continuity $\mathbb{P}^{h}_{T}\ll\mathbb{P}^{g}_{T}$ again follows from 3.6.
@@ 4.2. Choosing $g$ using the transition density of a scaled Brownian motion
By C.1, $g$ in Equation 4.8 is bounded from above and bounded away from 00. Since $g$ is smooth in $t$ and well-defined in $T$, we clearly have that $g\in{\mathcal{D}}({\mathcal{A}})$ and continuous differentiability of the maps $t\mapsto g(t,x)$ for all $x\in{\mathbb{S}}$ on $[0,T]$. The result thus follows from 3.6. ∎
@@
$M(t):=\left(C+L{a}L^{\prime}(T-t)\right)^{-1},$
then $H(t)=L^{\prime}M(t)L$ and $F(t)=L^{\prime}M(t)v$ and therefore
The proportionality is up to a differentiable time-dependent function and hence does not affect the change of measure by 3.3. The result now follows by substituting this expression into the definition of $\alpha^{g}_{\ell}(t,x)$ and rewriting in terms of the metric $d$. ∎
This implies that only one matrix inverse, $(LaL^{\prime})^{-1}$, needs to be calculated. By Equation 4.6, the intensity of the guided process either becomes very large or small depending on whether $L(x+\xi_{\ell})$ is closer or further away from $v$ than $Lx$ with respect to the metric $d$, which has intuitive appeal. Also note that $\varepsilon$ appears only in the denominator at $T+\varepsilon$, which implies that the choice $C=\varepsilon L{a}L^{\prime}$ conditions the process at time $T+\varepsilon$, thereby precluding explosive behaviour in (4.8) as $t\uparrow T$.
@@
Ideally, we would like to have ${\mathbb{P}}^{g}\left(LX(T)=v\right)=1$. For $g=g_{\varepsilon}$ as in 4.4 this is not guaranteed. Essentially, the choice $C=\varepsilon L{a}L^{\prime}$ yields a process where $LX(t)$ hits $v$ at time $T+\varepsilon$. Since the intensities are bounded from below and above, there is a positive probability of at least one reaction occurring in $(T,T+\varepsilon)$ resulting in $LX(T)\neq v$. In Section 4.3 we discuss the case where $\varepsilon=0$, which does yield ${\mathbb{P}}^{g}(LX(T)=v)=1$.
@@
Comparing $g_{\mathrm{F}}$, $g_{\mathrm{LNAR}}$ and $g_{\varepsilon}$, it is conceivable that $g_{\varepsilon}$ will deviate most from $h$. From a computational point of view $g_{\varepsilon}$ is most attractive, as can be seen from (4.6). Moreover, in Section 5 it turns out that simulation of the guided process is simplified in case of $g_{\varepsilon}$.
@@
In dimension $1$, we derive a better intuition for the choice for $a$ from (4.6) and (4.7). A small choice leads to a higher guided intensity, and thus a process arriving around $v$ quickly and likely to stay near $v$, while a large value of $a$ leads to trajectories that remain unaffected by $\alpha_{\ell}^{g_{\varepsilon}}$ until $t$ approaches $T$. Ideally, we mimic trajectories under ${\mathbb{P}}^{h}$ and thus a good choice of $a$ ensures $g(0,x_{0})\Psi_{T}(X)\mathbf{1}\{LX(T)=v\}/g(T,X(T-))$ is large (with $\Psi_{t}(X)$ defined in Equation (3.8)). This can be verified using Monte-Carlo simulation sampling $X$ under ${\mathbb{P}}^{g_{\varepsilon}}$. Alternatively, a simple choice for $a$ is $a_{\mathrm{CLE}}(0,x_{0})$ or, if we have a complete observation, $a=a_{\mathrm{CLE}}(T,v)$.
@@ 4.2.1. Extension to multiple observations
For each observation $k$, we consider $V_{k}\mid X(t_{k})\sim{\mathcal{N}}(0,C_{k})$ where $C_{k}$ is an $m_{k}\times m_{k}$ covariance matrix. Suppose $q_{k}$ denote the density of the ${\mathcal{N}}(0,C_{k})$-distribution. Then the transition density of $\tilde{X}$ satisfies for $t\in[t_{k-1},t_{k})$,
@@
Using C.2, the proof of 4.1 can be repeated for a result like 4.1 for each observation $k$ by taking limits $t\uparrow t_{k}$ to find that ${\mathbb{P}}_{t_{n}}^{h}\ll{\mathbb{P}}_{t_{n}}^{g}$ with
@@
4.3. Choosing $g$ so that ${\mathbb{P}}^{h}$ and ${\mathbb{P}}^{g}$ are equivalent
@@
By 4.5, the choice $C=\varepsilon LaL^{\prime}$ can be interpreted as imposing a condition of hiting $v$ at time $T+\varepsilon$. Therefore, intuitively, equivalence can be achieved upon artificially setting all elements in the matrices $C_{k}$ equal to 00. In this case, $g(t_{k},x)$ becomes ill-defined whenever $L_{k}x\neq v_{k}$ and boundedness of $g$ is lost, rendering the earlier proofs invalid. To utilize this choice for $g$ and show equivalence, we build on earlier work in Corstanje et al., (2023).
4.3.1. Setting and absolute continuity
@@
where $L$, $v$ and $M$ are defined in 4.9 but with $C_{1},\dots,C_{n}=0$ and we define $g(t_{k},x)=g(t_{k}+,x)$ for $x\in L_{k}^{-1}v_{k}$. We now formalize one particular property of $g$ that is important for the remainder of the session.
Lemma 4.10.
$\lim_{t\uparrow t_{k}}g(t,x)=\begin{cases}g(t_{k},x)&\text{if }L_{k}x=v_{k}\\ 0&\text{otherwise}\end{cases}.$ (4.16)
Proof.
$(v(t)-L(t)x)^{\prime}M(t)(v(t)-L(t)x)=\frac{d_{k}(v_{k},L_{k}x)^{2}}{(t_{k}-t)}+o(t_{k}-t),$ (4.17)
where $d_{k}$ is the metric on ${\mathbb{R}}^{m_{k}}$ defined as
$d_{k}(x,y)=\sqrt{(y-x)^{\prime}\left(L_{k}a_{k}L_{k}^{\prime}\right)^{-1}(y-x)},\qquad x,y\in{\mathbb{R}}^{m_{k}}.$ (4.18)
which yields (4.16). Since $v$ and $L$ are piecewise constant and $\frac{\,\mathrm{d}M}{\,\mathrm{d}t}=M(t)\begin{pmatrix}L_{i}a_{k}L_{j}^{\prime}\end{pmatrix}_{i,j=k}^{n}M(t)$ is positive semidefinite, $\partial_{t}\log g(t,x)\leq 0$. ∎
@@
Proof.
The proof is deferred to Appendix B. ∎
To show that ${\mathbb{P}}^{g}_{t_{n}}(A[X^{t_{n}}])>0$, we require 4.12, combined with the observation that the proof of 4.2 can be repeated with this choice for $g$ upon observing that $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}$ stays finite on the sets $\{L_{k}x=v_{k}\}$
@@
$A_{n}\subseteq A[X^{t_{n}}]$
Proof.
The proof is deferred to Appendix B
@@
4.3.2. Equivalence
(4.19) shows that ${\mathbb{P}}^{h}$ is absolutely continuous with respect to ${\mathbb{P}}^{g}$ on the set $A[X^{t_{n}}]$. For simulation purposes, equivalence would be preferable. It follows from (4.19) that this is indeed the case if ${\mathbb{P}}^{g}\left(A[X^{t_{n}}]\right)=1$. This can be shown if the network also satisfies a greedy property.
@@ 4.3.2. Equivalence
For $k=1,\dots,n$, define the metric $d_{k}$ on ${\mathbb{R}}^{m_{k}}$ through (4.18) and suppose that $d_{1},\dots,d_{n}$ are such that for all $k$, $t\in[t_{k-1},t_{k})$ $x\in{\mathbb{S}}\setminus L_{k}^{-1}v_{k}$, there is a reaction $\ell\in{\mathcal{R}}$ such that $\lambda_{\ell}(t,x)>0$ and $d_{k}(v_{k},L_{k}(x+\xi_{\ell}))<d_{k}(v_{k},L_{k}x)$.
Intuitively, 4.14 is satisfied when there is always a reaction available that takes the process closer to the first desired conditioning given being in state $x$ at time $t$.
@@
Suppose 4.14 is satisfied. Then ${\mathbb{P}}^{g}_{t_{n}}(A_{n})=1$.
Proof.
The proof is deferred to Appendix B
@@
It follows from 4.12 and 4.15 that, under 4.14, ${\mathbb{P}}^{g}_{t_{n}}(A[X^{t_{n}}])=1$.
@@
4.4. Choosing $g$ to guide a process with monotone components
4.1 and 4.11 prove that choosing $g$ using the transition density of a scaled Brownian motion yields a process that is absolutely continuous with respect to the conditioned process. However, depending on the network, this might not always be a desirable choice.
Consider for instance 2.5. Here, the fourth component of the process (P) is monotonically increasing as it only appears in reaction $3$, where $1$ is added, which can also be seen in Figure 2. Let us for simplicity assume a complete observation at time $T$, that is $X(T)=v_{T}$. Clearly, if $X_{4}(t)=v_{T,4}$ for some $t<T$, reaction $3$ cannot occur anymore for the process to satisfy the conditioning. However, $\lambda_{3}^{g}(t,X(t))\neq 0$ for the choices of $g$ discussed so far. This can lead to trajectories that don’t satisfy the conditioning, which in turn means a low acceptance ratio when sampling.
$g(t,(z,y))=\exp\left(-\frac{d(z_{T},z)^{2}}{2(T-t)}\right)\frac{\left(\tilde{\theta}(T-t)\right)^{y_{T}-y}}{(y_{T}-y)!}\exp\left(-\tilde{\theta}(T-t)\right),$ (4.20)
where
$d(z_{T},z)=\sqrt{(z_{T}-z)^{\prime}a^{-1}(z_{T}-z)}$
Clearly,
$\lim_{t\uparrow T}g(t,(z,y))=\begin{cases}1&\text{if }(z,y)=(z_{T},y_{T})\\ 0&\text{otherwise }.\end{cases}$
Since also $\partial_{t}\log g\leq 0$, 4.10 is also holds for this choice of $g$, meaning 4.11 still applies. However, now $\lambda_{3}^{g}(t,x)=0$ whenever $x_{4}=v_{T,4}$.
@@ 5. Simulation methods
In this section we discuss methods for simulating the guided process on $[0,T]$. This is a Markov jump process with time-dependent intensity which can become large when time is close to $T$. Simulating a Markov jump process is easy when intensities do not depend on time. In this case, given a state $x$ at time $t$, we simply simulate reaction times $\tau_{\ell}\sim\mathrm{Exp}(\lambda_{\ell}(x))$ for each reaction, set ^$\hat{\ell}=\operatornamewithlimits{argmin}_{\ell\in{\mathcal{R}}}\tau_{\ell}$ and move $t\leftarrow t+\tau_{\hat{\ell}}$ and $X(t+\tau_{\hat{\ell}})\leftarrow X(t)+\xi_{\hat{\ell}}$. When ${\mathcal{R}}$ contains many reactions, one could alternatively use Gillespie’s algorithm, see e.g. Gillespie, (1976, 1977), which only samples the first reaction time and then searches for the most likely reaction at that time.
To extend this method to chemical reaction processes with time-dependent intensities, we have to sample reaction times whose distribution satisfies (2.4), for which there is generally no direct sampling method. We therefore apply a Poisson thinning step, which results in in Algorithm 1.

### v2

@@ 4. Choices for $g$
It remains to specify the maps $(t,x)\mapsto g(t,x)$ satisfying the assumption of 3.3. Moreover, to be of any use, the event $\{LX_{T}=v\}$ needs to get positive probability under ${\mathbb{P}}^{g}_{T}$.
In this section we show the following:
  • Section 4.1: The choices for $g$ in Fearnhead, (2008) and Golightly & Sherlock, (2019), both based on the chemical Langevin equation, lead to absolute continuity.
  • Section 4.2: A guiding function based on the transition density of a scaled Brownian motion yields absolute continuity while being computationally very efficient. It satisfies ${\mathbb{P}}^{g}_{T}(LX(T)=v)>0$. In subsection 4.2.1 we show how $g$ can be extended to the case of multiple partial observations.
  • Section 4.3: For processes with monotone components a specific choice for $g$ can overcome certain problems arising in the aforementioned choices.
In the following, $C$ denotes a positive definite matrix.
@@ 4.1. Choices based on the chemical Langevin equation
$g_{\rm F}(t,x)={\mathcal{N}}\left(v;L(x+b_{\mathrm{CLE}}(t,x)(T-t)),La_{\mathrm{CLE}}(t,x)L^{\prime}(T-t)+C\right),$ (4.1)
As the assumption of 3.5 is satisfied we obtain ${\mathbb{P}}^{h}_{T}\ll{\mathbb{P}}^{g_{\mathrm{F}}}_{T}$.
Golightly & Sherlock, (2019) propose to infer $g$ by using the Linear Noise Approximation (LNA) to the CLE. Here, we discuss their specific choice called “LNA with restart”. For notational convenience, write $b=b_{\mathrm{CLE}}$ and $\sigma=\sigma_{\mathrm{CLE}}$ and again set $a=\sigma\sigma^{\prime}$. Given $t<T$ and $x\in{\mathbb{S}}$, we denote by $z_{T\mid(t,x)}$ and $V_{T\mid(t,x)}$ the solutions at time $T$ of the system of ordinary differential equations
@@
Absolute continuity $\mathbb{P}^{h}_{T}\ll\mathbb{P}^{g_{\mathrm{LNAR}}}_{T}$ follows by the same argument used for $g_{\mathrm{F}}$.
@@ 4.2. Choosing $g$ using the transition density of a scaled Brownian motion
$M(t):=\left(C+L{a}L^{\prime}(T-t)\right)^{-1},$
then $H(t)=L^{\prime}M(t)L$ and $F(t)=L^{\prime}M(t)v$ and therefore
The proportionality is up to a differentiable time-dependent function and hence does not affect the change of measure by 3.7. By C.1, $g$ is bounded from above and bounded away from 00. Since $g$ is smooth in $t$ and well-defined in $T$, we clearly have continuous differentiability of the maps $t\mapsto g(t,x)$ for all $x\in{\mathbb{S}}$ on $[0,T]$. The result thus follows from 3.5 and 3.4. ∎
@@
This result follows by substituting the expression for $g$ in (4.7) into the definition of $\alpha^{g}_{\ell}(t,x)$ and rewriting in terms of the metric $d$. ∎
This implies that only one matrix inverse, $(LaL^{\prime})^{-1}$, needs to be calculated. By Equation 4.8, the intensity of the guided process either becomes very large or small depending on whether $L(x+\xi_{\ell})$ is closer or further away from $v$ than $Lx$ with respect to the metric $d$, which has intuitive appeal. Also note that $\varepsilon$ appears only in the denominator at $T+\varepsilon$, which implies that the choice $C=\varepsilon L{a}L^{\prime}$ imposes the conditioning at at time $T+\varepsilon$ instead of $T$, thereby precluding explosive behaviour in (4.7) as $t\uparrow T$.
@@
Ideally, we would like to have ${\mathbb{P}}^{g}\left(LX(T)=v\right)=1$. For $g=g_{\varepsilon}$ as in 4.4 this is not guaranteed. Essentially, the choice $C=\varepsilon L{a}L^{\prime}$ yields a process where $LX(t)$ hits $v$ at time $T+\varepsilon$. Since the intensities are bounded from below and above, there is a positive probability of at least one reaction occurring in $(T,T+\varepsilon)$ resulting in $LX(T)\neq v$. In Section 5 we discuss the case where $\varepsilon=0$, which does yield ${\mathbb{P}}^{g}(LX(T)=v)=1$.
@@
Comparing $g_{\mathrm{F}}$, $g_{\mathrm{LNAR}}$ and $g_{\varepsilon}$, it is likely that $g_{\varepsilon}$ will deviate most from $h$. In particular when the drift and diffusion coefficient of the CLE are nonlinear, $g_{\varepsilon}$ will likely deviate from $h$ the most leading to the guided paths deviating from the conditioned paths. However, from a computational point of view $g_{\varepsilon}$ is most attractive, as can be seen from (4.8). Moreover, in Section 6 it turns out that simulation of the guided process is simplified in case of $g_{\varepsilon}$.
@@
In dimension $1$, we derive a better intuition for the choice for $a$ from (4.8) and (4.9). A small choice leads to a higher guided intensity, and thus a process arriving around $v$ quickly and likely to stay near $v$, while a large value of $a$ leads to trajectories that remain unaffected by $\alpha_{\ell}^{g_{\varepsilon}}$ until $t$ approaches $T$. Ideally, we mimic trajectories under ${\mathbb{P}}^{h}$ and thus a good choice of $a$ ensures $g(0,x_{0})\Psi_{T}^{g}(X)\mathbf{1}\{LX(T)=v\}/g(T,X(T))$ is large (with $\Psi_{t}^{g}(X)$ defined in Equation (3.8)). This can be verified using Monte-Carlo simulation sampling $X$ under ${\mathbb{P}}^{g_{\varepsilon}}$. Alternatively, a simple choice for $a$ is $a_{\mathrm{CLE}}(0,x_{0})$ or, if we have a complete observation, $a=a_{\mathrm{CLE}}(T,v)$.
@@ 4.2.1. Extension to multiple observations
For each observation $k$, we consider $V_{k}\mid X(t_{k})\sim{\mathcal{N}}(0,C_{k})$ where $C_{k}$ is an $m_{k}\times m_{k}$ covariance matrix. Suppose $q_{k}$ denotes the density of the ${\mathcal{N}}(0,C_{k})$-distribution. Then the transition density of $\tilde{X}$ satisfies for $t\in[t_{k-1},t_{k})$,
@@
Using C.2, the proof of 4.1 can be repeated for a result like 4.1 for each observation $k$ to find that ${\mathbb{P}}_{t_{n}}^{h}\ll{\mathbb{P}}_{t_{n}}^{g}$ with
@@
4.3. Choosing $g$ to guide a process with monotone components
4.1 shows that choosing $g$ using the transition density of a scaled Brownian motion yields a process that is absolutely continuous with respect to the conditioned process. However, depending on the network, this might not always be a desirable choice.
Consider for instance 2.4. Here, the fourth component of the process (P) is monotonically increasing as it only appears in reaction $3$, where $1$ is added, which can also be seen in Figure 2. Let us for simplicity assume a complete observation at time $T$, that is $X(T)=v_{T}$. Clearly, if $X_{4}(t)=v_{T,4}$ for some $t<T$, reaction $3$ cannot occur anymore for the process to satisfy the conditioning. However, $\lambda_{3}^{g}(t,X(t))\neq 0$ for the choices of $g$ discussed so far. This can lead to trajectories that don’t satisfy the conditioning, which in turn means a low acceptance ratio when sampling.
$g(t,(z,y))=\exp\left(-\frac{d(z_{T},z)^{2}}{2(T+\epsilon-t)}\right)\frac{\left(\tilde{\theta}(T-t)\right)^{y_{T}-y}}{(y_{T}-y)!}\exp\left(-\tilde{\theta}(T-t)\right),$ (4.16)
where $\epsilon>0$ and
$d(z_{T},z)=\sqrt{(z_{T}-z)^{\prime}a^{-1}(z_{T}-z)}.$
This choice satisfies the assumption of 3.3.
5. Choosing $g$ so that ${\mathbb{P}}^{h}$ and ${\mathbb{P}}^{g}$ are equivalent
@@
By 4.5, the choice $C=\varepsilon LaL^{\prime}$ can be interpreted as imposing a condition of hiting $v$ at time $T+\varepsilon$. Therefore, intuitively, equivalence can be achieved upon artificially setting all elements in the matrices $C_{k}$ equal to 00. In this case, $g(t_{k},x)$ becomes ill-defined whenever $L_{k}x\neq v_{k}$ and boundedness of $g$ is lost, rendering the earlier proofs invalid. To utilize this choice for $g$ and show equivalence, we build on earlier work in Corstanje et al., (2023). Proofs of results in this section are given in Appendix B.
5.1. Absolute continuity
@@
where $L$, $v$ and $M$ are defined in 4.9 but with $C_{1},\dots,C_{n}=0$ and we define $g(t_{k},x)=g(t_{k}+,x)$ for $x\in L_{k}^{-1}v_{k}$.
@@
To show that ${\mathbb{P}}^{g}_{t_{n}}(A[X^{t_{n}}])>0$, we require 5.2, combined with the observation that the proof of 4.2 can be repeated with this choice for $g$ upon observing that $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}^{g}$ stays finite on the sets $\{L_{k}x=v_{k}\}$.
@@
$A_{n}\subseteq A[X^{t_{n}}]$ with $A_{n}$ as defined in (1.2).
@@
5.2. Equivalence
5.1 shows that ${\mathbb{P}}^{h}$ is absolutely continuous with respect to ${\mathbb{P}}^{g}$ on the set $A[X^{t_{n}}]$. For simulation purposes, equivalence would be preferable. It follows from (5.2) that this is indeed the case if ${\mathbb{P}}^{g}\left(A[X^{t_{n}}]\right)=1$. This can be shown if the network also satisfies a greedy property.
@@ 5.2. Equivalence
For $k=1,\dots,n$, define the metric $d_{k}$ on ${\mathbb{R}}^{m_{k}}$ through (B.6) and suppose that $d_{1},\dots,d_{n}$ are such that for all $k$, $t\in[t_{k-1},t_{k})$ $x\in{\mathbb{S}}\setminus L_{k}^{-1}v_{k}$, there is a reaction $\ell\in{\mathcal{R}}$ such that $\lambda_{\ell}(t,x)>0$ and $d_{k}(v_{k},L_{k}(x+\xi_{\ell}))<d_{k}(v_{k},L_{k}x)$.
Intuitively, 5.4 is satisfied when there is always a reaction available that takes the process closer to the first desired conditioning given being in state $x$ at time $t$. The choice for $g$ will then guarantee that eventually, the path will jump to this point closer to the conditioned state and will therefore hit the observation in finitely many reactions. This argument is formalized in 5.5.
@@
Suppose 5.4 is satisfied. Then ${\mathbb{P}}^{g}_{t_{n}}(A_{n})=1$.
@@
It follows from 5.2 and 5.5 that, under 5.4, ${\mathbb{P}}^{g}_{t_{n}}(A[X^{t_{n}}])=1$.
@@ 6. Simulation methods
In this section we discuss methods for simulating the guided process on $[0,T]$. This is a Markov jump process with time-dependent intensity. In case $\|LX_{t}-v\|$ is large for $t$ close to $T$, this intensity may blow up.
Simulating a Markov jump process is easy when intensities do not depend on time. In this case, given a state $x$ at time $t$, we simply simulate reaction times $\tau_{\ell}\sim\mathrm{Exp}(\lambda_{\ell}(x))$ for each reaction, set ^$\hat{\ell}=\operatornamewithlimits{argmin}_{\ell\in{\mathcal{R}}}\tau_{\ell}$ and move $t\leftarrow t+\tau_{\hat{\ell}}$ and $X(t+\tau_{\hat{\ell}})\leftarrow X(t)+\xi_{\hat{\ell}}$. When ${\mathcal{R}}$ contains many reactions, one could alternatively use Gillespie’s algorithm, see e.g. Gillespie, (1976, 1977), which first samples the reaction time and subsequently the reaction that takes place at that time.
To extend this method to chemical reaction processes with time-dependent intensities, we have to sample reaction times satisfying (2.4). In general this is hard and therefore therefore we consider a Poisson thinning step. This gives Algorithm 1. The efficiency of Algorithm 1 will depend on whether sharp bounds $\bar{\lambda}_{\ell}$ can be derived.